OpenStudy (anonymous):

please help solve the system using substitution put your answer in ordered pair (x,y) 8x-y=4 5x+3y=17

OpenStudy (anonymous):

ok so look at one of the equations, lets use the top one 8x - y = 4 can you solve this for y? it will be in terms of x

OpenStudy (anonymous):

@priea

OpenStudy (anonymous):

im a little confused....sorry @billj5

OpenStudy (anonymous):

can you solve 8x - y = 4 for x?

OpenStudy (anonymous):

let me put it another way if i give you 8x - y = 4 what does x equal?

OpenStudy (anonymous):

4x-y?

OpenStudy (anonymous):

ok when you are trying to find out what a variable equals in an equation, you need to get that variable by itself on one side of the equals sign (one side of the equation)

OpenStudy (anonymous):

lets try an easier one

OpenStudy (anonymous):

x + 5 = 10 what does x equal?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

correct you just got x by itself by subtracting 5 from both sides

OpenStudy (anonymous):

ok now try again with 8x - y = 4

OpenStudy (anonymous):

what is the first thing you would do to get x by itself?

OpenStudy (anonymous):

ill give you a hint, you would add y to both sides

OpenStudy (anonymous):

8x - y + y = 4 + y

OpenStudy (anonymous):

-y + y = 0 so we have 8x + 0 = 4 + y which is 8x = 4 + y

OpenStudy (anonymous):

are you good with that part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok now what would you do to get x by itself?

OpenStudy (anonymous):

8x = 4 + y, get x by itself, somehow remove the 8 coefficient

OpenStudy (anonymous):

subtract 4 on both sides?

OpenStudy (anonymous):

what operation is the inverse of multiplication?

OpenStudy (anonymous):

ok if I have 8x and divide that by 8 what do i get?